Newsome Inc. buys on terms of 3/15, net 45. It does not take the discount, and it generally pays after 60 days. What is the nominal annual percentage cost of its non-free trade credit, based on a 365-day year?
Answer #1

3/15 Net 45 means a disc of 3% if paid within 15 days else make full payment within 45 days
So if a customer doesn't take discount, he gets free credit for 45-15 = 30 days
But here payment is coming after 60 days. ie 60-15 = 45 days free credit is being availed.
You are getting a 3% discount on a 45-day advance in payments.
There are (365/45) = 8.11 45-day periods in a year.

Nominal cost of trade credit = (Disc %age/100-Disc %age)*[365/(Days credit o/s - Disc period)]

ie Nominal cost of trade credit = (3%/97%)*(365/(60-15)) = 3.09%* 8.11 = 25.09%

EAR = (1+3.09%)^8.11 - 1 = 27.99%.